Suppose you intend to estimate ln(1.2) using a linearization ?(?) = ?(?) + ?′(?)(? − ?). a. What are appropriate choices for ?(?) and ?? b. Set up and evaluate the computation that would give you the desired estimate for ln(1.2).
Answer #1

L(n) = f(a) + f(a) (x-a). To estimate ls (1.2) f(x) = ln (w.) and a = L(x) = f(1) + f(& (n-1). f(x) = ln (2) - ln (1) + (x-